OpenStudy (anonymous):

What is the slope of the line 3x + 2y =5? (MEDAL FOR BEST ANSWER!)

OpenStudy (anonymous):

2y = -3x + 5 y = -3/2x + 5 Slope = -3/2

OpenStudy (anonymous):

INDEED

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

No problem :) Any time

OpenStudy (dan815):

hi